OpenStudy (anonymous):

What's the answer of "sin x =4/15?"

OpenStudy (asevilla5):

x = sin-1 (4/15) ( sin -1 is [SHIFT] [sin] on a calculator ) = 15.5

OpenStudy (mathmale):

Don't you mean, "What is the solution of the equation sin x = 4/15 ?" You posted a similar problem a few minutes ago. What did you learn from working on it?

OpenStudy (mathmale):

Please note that sin x = 4/15 has multiple solutions. Were you supposed to find all of the solutions, or just the solutions on a given interval?

OpenStudy (anonymous):

Accidentally had a typo sorry, and thanks. :)

OpenStudy (anonymous):

This mistake won't happen again, i promise you.

OpenStudy (mathmale):

for example, sin x = 4/15 has two positive roots on the interval [0,pi] and two negative roots on the interval [pi, 2pi].